OpenStudy (anonymous):

which is the solution to the equation 3x -4=5(x-4)?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you have so far

jimthompson5910 (jim_thompson5910):

absolute value? I don't see any absolute value bars in the original problem `3x -4=5(x-4)`

OpenStudy (anonymous):

Mistake on my part.

jimthompson5910 (jim_thompson5910):

can you draw out the full problem?

OpenStudy (anonymous):

\[3x-4=5(x-4)\]

OpenStudy (anonymous):

like that?

jimthompson5910 (jim_thompson5910):

are there any vertical bars in the problem?

OpenStudy (anonymous):

I made a big mistake. Hold on.

OpenStudy (anonymous):

x=0?

OpenStudy (anonymous):

X=0 because different multiplications of X are being subtracted by 4. 3x3=9-4=5 5x3=15-4=9 5 does not equal 9. 3x0=0-4=-4 and 5x0=0-4=-4 but negative 4 equals negative 4 so the answer is X= 0?

jimthompson5910 (jim_thompson5910):

is it possible to post a screenshot? that would clear up all confusion

OpenStudy (anonymous):

The question isn't online.

jimthompson5910 (jim_thompson5910):

oh it's on paper, I see

jimthompson5910 (jim_thompson5910):

I'm going to assume there is no set of absolute value bars The first step is to distribute the 5 through 3x -4=5(x-4) 3x -4=5*x-5*4 3x -4 = 5x - 20 with me so far?

OpenStudy (anonymous):

There are no bars. I am with you

jimthompson5910 (jim_thompson5910):

ok now we need to move the x terms over to one side. Say the left side

jimthompson5910 (jim_thompson5910):